Determine the energy in joules and electron volts of a quantum with a wavelength of 600.0nm


1
Expert's answer
2021-04-19T17:07:26-0400

The energy of quantum is given as follows:


"E = h\nu"

where "h \approx 6.626\times 10^{-34}J\cdot s" or "h = 4.136\times 10^{-15}eV\cdot s" is the Plank's constant, and

"\nu = \dfrac{c}{\lambda}"

is the frequency of the quantum, where "c = 3\times10^8m/s" is the speed of light, and "\lambda = 600nm = 6\times 10^{-7}m" is the wavelength.

Thus, obtain for joules:


"E = \dfrac{hc}{\lambda} = \dfrac{6.626\times 10^{-34}\cdot 3\times 10^8}{6\times 10^{-7}} \approx 3.313\times10^{-19}J"

In electron volts:


"E = \dfrac{hc}{\lambda} = \dfrac{4.136\times 10^{-15}\cdot 3\times 10^8}{6\times 10^{-7}} \approx 2.068eV"

Answer. "3.313\times10^{-19}J" or "2.068eV".